Find the volume V of the solid obtained by rotating the region bounded by the given curves about the specified line. y = 5(sqt 25 − x2) , y = 0, x = 0, x = 3; about the x-axis
Answer #1

The regean boundled by - 5 125-12 ll v=8eyhde = 23 25 (25x3 de = 255 [254-2373 = 250 [25x3-9-o - 250 x 66 = 165085 V- 1650T